### v1

# DeSTA2.5-Audio: Toward General-Purpose Large Audio Language Model with Self-Generated Cross-Modal Alignment

Abstract
We introduce DeSTA2.5-Audio, a general-purpose Large Audio Language Model (LALM) designed for robust auditory perception and instruction-following, without requiring task-specific audio instruction-tuning. Recent LALMs typically augment Large Language Models (LLMs) with auditory capabilities by training on large-scale, manually curated or LLM-synthesized audio-instruction datasets. However, these approaches have often suffered from the catastrophic forgetting of the LLM’s original language abilities. To address this, we revisit the data construction pipeline and propose DeSTA, a self-generated cross-modal alignment strategy in which the backbone LLM generates its own training targets. This approach preserves the LLM’s native language proficiency while establishing effective audio-text alignment, thereby enabling zero-shot generalization without task-specific tuning. Using DeSTA, we construct DeSTA-AQA5M, a large-scale, task-agnostic dataset containing 5 million training samples derived from 7,000 hours of audio spanning 50 diverse datasets, including speech, environmental sounds, and music. DeSTA2.5-Audio achieves state-of-the-art or competitive performance across a wide range of audio-language benchmarks, including Dynamic-SUPERB, MMAU, SAKURA, Speech-IFEval, and VoiceBench. Comprehensive comparative studies demonstrate that our self-generated strategy outperforms widely adopted data construction and training strategies in both auditory perception and instruction-following capabilities. Our findings underscore the importance of carefully designed data construction in LALM development and offer practical insights for building robust, general-purpose LALMs.  [^1] https://github.com/kehanlu/DeSTA2.5-Audio
Index Terms:
Large Audio Language Model, cross-modal alignment, dataset construction, instruction-tuning
I Introduction
The development of general-purpose artificial intelligence has become a central focus in contemporary AI research, driven by the remarkable performance of large language models (LLMs) across various natural language understanding and generation tasks [1, 2, 3, 4, 5, 6, 7]. Building on these advancements, a promising direction is to equip LLMs with multi-modal understanding capabilities, leading to the emergence of Large Audio Language Models (LALMs) [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] and Large Vision Language Models (LVLMs)[23, 24, 25, 26, 27]. This paper focuses on building a general-purpose LALM, illustrated in Figure 1.
To develop a general-purpose LALM, two core capabilities are essential: auditory perception and instruction-following. Auditory perception refers to the comprehensive processing of auditory information, including speech, non-verbal cues, background sounds, and music. In addition, instruction-following entails interpreting diverse user commands, integrating them with internal knowledge, and producing contextually appropriate responses. To achieve this, most LALM architectures adopt a modular approach, combining a pre-trained audio model with a text-based LLM as the knowledge backbone [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. These components are integrated through a cross-modal alignment process, which typically involves joint fine-tuning on a corpus of paired audio and text instructions to bridge the representational gap between the audio and text modalities. Accordingly, the research problem centers on training the fusion model with limited audio-related data, while leveraging the LLM’s generalization capabilities to handle unseen tasks.
Figure 1: A general-purpose Large Audio Language Model (LALM) understands multifaceted audio information and accepts arbitrary text prompts to generate corresponding responses, enabling seamless multimodal interaction.
However, acquiring high-quality data for cross-modal alignment remains a significant challenge. A common strategy involves transforming existing audio datasets by converting them into audio instruction-tuning datasets, consisting of (1) audio input, (2) textual instructions, and (3) textual responses. These training triplets can be generated either by text-based LLMs or through manual annotation. To improve the generalization capabilities of LALMs, several techniques have been adopted, including fine-tuning lightweight LoRA adapters[11, 10, 12, 14, 13], adopting multi-stage curriculum learning [12, 16, 17, 13, 14, 15, 20], and applying test-time pruning [12]. While current LALMs perform well across a range of audio-related benchmarks [28, 29, 30, 31, 32, 33, 34], many studies highlight the issue of catastrophic forgetting [8, 9, 12, 11, 10, 14, 35, 36], where the models become overly specialized to a limited set of audio-related tasks at the cost of generalization to unseen tasks. Moreover, our previous work, Speech-IFEval [36], also reveals that most LALMs experience significant degradation in text-based knowledge, underscoring the persistent limitations of current cross-modal alignment techniques.
To address the issue of catastrophic forgetting in cross-modal alignment, we revisit the data construction process with a key observation: annotations from different sources, such as various LLMs or human annotators, exhibit distinct stylistic and behavioral patterns. While this may seem intuitive, prior studies often rely on responses generated by powerful LLMs [11, 10, 16, 17, 13], such as GPT [1], or human-designed pairs [12, 14] to ensure high-quality supervision. However, we hypothesize that incorporating annotations from heterogeneous sources inevitably introduces biases into the dataset. As a result, the fusion model must not only learn auditory grounding but also implicitly adapt to the new response patterns introduced by the annotators. We believe this extra adaptation burden can gradually distort the LLM’s native output distribution or instruction-following behavior, ultimately degrading its text-based knowledge.
Figure 2: (Left) Dataset Construction: An audio description $x^{\text{text}}$ and a randomly sampled prompt $\mathbf{p}$ are fed into the backbone LLM to generate training targets $\mathbf{y}$. (Right) Model Training: The fusion model is trained using the self-generated targets $\mathbf{y}$ along with the corresponding audio inputs $x^{\text{audio}}$ and prompt $\mathbf{p}$. The fire and snowflake icons indicate trainable and frozen modules, respectively. The audio decoder is optional.
Building on this insight, we introduce DeSTA [^2] DeSTA stands for Descriptive Speech-Text Alignment, first introduced in [8] and further improved with self-generated design in [9]. While the method generalizes beyond speech in this study, we retain the original naming convention for consistency., a self-generated cross-modal alignment framework that mitigates conflicting supervision by having the backbone language model generate its own training targets. Specifically, we convert the metadata of each audio segment into a structured textual description and pair it with an arbitrary prompt. Then, the LLM generates a corresponding response, which serves as the training target for cross-modal alignment. This self-generated supervision ensures stylistic and semantic consistency with the LLM’s native output distribution, thereby preserving its instruction-following capabilities while enabling effective adaptation to auditory inputs. In our prior work, DeSTA2 [9], we demonstrate that training solely on a single general task, speech captioning, is sufficient for learning robust audio grounding. The resulting model exhibited strong zero-shot generalization across a wide range of downstream tasks without any task-specific instruction tuning, presenting a scalable path toward general-purpose LALM training. In this work, we extend DeSTA by scaling our dataset construction pipeline across diverse audio domains, including speech, environmental sounds, and music. We aggregate 50 publicly available audio datasets spanning a wide range of audio information to create a large-scale corpus of approximately 5 million audio-text pairs, which we call DeSTA-AQA5M. This dataset encompasses various forms of audio information, including paralinguistic cues, speaker identities, audio quality indicators, and environmental or contextual sounds. This expanded model, referred to as DeSTA2.5-Audio, further validates the effectiveness of self-generated alignment in enabling general and robust audio-language modeling.
We evaluate DeSTA2.5-Audio on several audio-language benchmarks, including Dynamic-SUPERB [33, 32], MMAU [28], SAKURA [29], Speech-IFEval [36], and VoiceBench [30]. Despite using only 7,000 hours of training data, compared to 510,000 hours used by the baseline Qwen2-Audio-Instruct [17], our model achieves state-of-the-art or competitive results across domains. Our comparative analysis also emphasizes the critical role of the distribution mismatch problem. Models trained on self-generated targets consistently outperform those trained on mismatched data sources (e.g., from other LLM), even when the latter originate from larger or more powerful LLMs. Although adaptation techniques like LoRA [37] can alleviate distribution mismatches, they often come at the cost of reduced performance in complex reasoning or out-of-domain scenarios. These results suggest that training data distribution should be a fundamental design consideration in the development of LALMs.
The contribution can be summarized as follows:
  • We propose DeSTA, a self-generated cross-modal alignment method that leverages the backbone LLM to generate its own training targets, eliminating the need for task-specific instruction-tuning. By reducing the mismatch between training data and the model’s original behavior, DeSTA prevents the catastrophic forgetting problem.
  • We introduce a large-scale audio instruction-tuning dataset, DeSTA-AQA5M, which covers a wide array of audio features and has 5 million audio-prompt-response triplets. The dataset comprises 7,000 hours of diverse audio data spanning speech, environmental sounds, and music from 50 publicly available audio datasets.
  • Our model, DeSTA2.5-Audio, achieved superior results across multiple audio-language benchmarks. We demonstrate strong performance on both auditory perception and instruction-following tasks, showing significantly enhanced robustness compared to existing LALMs.
  • We conducted extensive comparison studies validating that self-generated data is critical for success, revealing significant limitations in existing data construction approaches. Our analysis demonstrates that using different LLMs for data generation leads to suboptimal performance or model collapse, and that even stronger models from the same family do not guarantee improvements, providing important insights and design considerations for future research.
II Related works
II-A Large Audio Language Model (LALM)
There are already several overview papers about LALMs; Therefore, we briefly review only the technologies directly relevant to this work. For a comprehensive overview, please refer to these overview papers [38, 39, 40, 41, 34].
Recent research on general-purpose LALMs has focused on addressing the text-audio modality gap by fine-tuning fusion modules with audio–text pairs [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Therefore, the availability and construction of training datasets play a critical role in model development. To efficiently create audio-text pairs, synthetic data generated by powerful text-based LLMs is widely adopted. This approach involves repurposing existing audio datasets and utilizing their annotated textual metadata to construct input-output pairs. For instance, LTU [11] prompts LLMs to generate open-ended audio question-answer pairs. Qwen-Audio [16] creates interactive data to enhance chat behavior. DeSTA [8] augments metadata into speech captions as a pre-training stage, enabling the model to learn multifaceted speech descriptions before the instruction-tuning phase. These synthetic processes can be easily automated and scaled. Once the initial textual metadata is collected, the training data can be generated accordingly. Consequently, scaling up audio datasets has emerged as another key area of research [42, 43, 44, 45]. Beyond audio-to-text models, an emerging research direction explores multi-modal LLMs designed for interactive scenarios that require both audio understanding and speech generation [46, 47, 48, 1, 49]. These models often build upon the architectural and training insights from LALMs, while placing greater emphasis on generating natural and context-aware speech responses.
II-B Self-Generation Mechanism
The concept of self-generated data has gained significant traction in text LLM fine-tuning [50, 51, 52]. Rather than relying on supervision from a powerful teacher model or human-annotated data, self-generated approaches enable the model to create its own training data. A key advantage of this method is that the training targets are naturally aligned with the model’s inherent behavior and data distribution. Several studies on text LLM have shown that this alignment can improve performance [53, 54, 55, 56, 57].
In multi-modal contexts, researchers have begun to adapt this approach to the speech domain. One common method uses automatic speech recognition (ASR) transcriptions as inputs and allows the backbone LLM to generate textual continuations as training targets [58, 59, 60, 61]. However, these efforts are often limited to spoken-content-centric tasks, such as ASR and spoken question answering, and tend to underutilize the rich, multidimensional information present in audio. Moreover, these studies rarely analyze how model behavior or performance might differ if the training data were not self-generated.
Our prior work, DeSTA2 [9], extended the self-generated paradigm in the speech domain by incorporating paralinguistic cues, speaker characteristics, and other non-verbal audio signals. We found that training on self-generated “speech captions,” which provide comprehensive audio grounding, enables the model to yield strong zero-shot performance on a variety of tasks, demonstrating the ability to generalize from a single training objective. In this work, we broaden our approach to encompass a wider variety of audio information, and we provide comprehensive evaluations and analyses to validate its effectiveness.
III Methodology
An overview of the DeSTA framework is shown in Figure 2. Just as humans can imagine auditory scenes from written narratives, such as voices and ambient sounds in a novel or film script, modern LLMs exhibit a similar abilities. Most auditory concepts, such as tone, speaking style, and environmental sounds, can be described in natural language. Based on this intuition, DeSTA aims to provide an approximate textual description of audio input, which will be used to generate the corresponding training target using the backbone LLM.
III-A Self-Generated Dataset Construction
During the data construction phase, we first collect a diverse set of datasets containing detailed metadata. The metadata associated with each audio segment is converted into a structured textual format following the schema below:
[timestamp] Spoken content (non-verbal attribute name: value)
For example, a speech clip can be represented as “[00:00-00:05] Hello world (Gender:Female, Emotion:Happy...)”, while an audio captioning sample could appear as “[00:00-00:10] (A dog barking)”. This conversion allows us to represent diverse audio types using a consistent textual format. The resulting initial dataset, denoted as $\mathcal{D}_{\text{initial}}=\{({x}^{\text{audio}},{x}^{\text{text}})\}$, aligns each audio clip ${x}^{\text{audio}}$ with its corresponding textual description ${x}^{\text{text}}$.
Based on these initial audio–description pairs, we utilize a text-based LLM to generate training targets that reflect how the model would naturally respond to the given descriptions. Specifically, we randomly sample a text prompt $\mathbf{p}$ from a predefined instruction pool $\mathcal{P}$, which includes diverse prompt types such as descriptive tasks (e.g., “Describe the audio”), role-playing scenarios (e.g., “Respond to the audio based on its expression”), and open-ended questions (e.g., “Where is the audio being recorded?”). These prompts are crafted to maximize the LLM’s ability to utilize all available information from the textual description, thereby enhancing the quality of cross-modal alignment. The backbone LLM then takes the text description $x^{\text{text}}$ and the prompt $\mathbf{p}$ as inputs and produces a response $\mathbf{y}=\text{LLM}(x^{\text{text}},\mathbf{p})$. This automated setup enables the generation of detailed and context-aware responses, eliminating the need for labor-intensive, task-specific instruction design. Through this procedure, we construct an audio instruction-tuning dataset in the form of $\mathcal{D}={(x^{\text{audio}},x^{\text{text}},\mathbf{p},\mathbf{y})}$, where each data point includes the audio input, its corresponding text description, the sampled prompt, and the LLM-generated response.
A key advantage of this method is that it preserves the LLM’s native interpretation and response behavior to the inputs, ensuring stylistic and semantic consistency in the training data. For example, we observe that the instruction-tuned Llama3.1 [4] typically generates explanatory responses, organizes content with bullet points, and often includes greeting phrases before the main content. These stylistic patterns, which are specific to the LLM, are naturally embedded in the generated training data. Thus, while the construction pipeline is compatible with any text-based LLM, employing the same model, i.e., the self-generated setting, is the most reasonable design for cross-modal alignment.
III-B Model Training
As illustrated in Figure 2 (Right), our model adopts a modular architecture that integrates a pre-trained audio model with an instruction-tuned LLM. To bridge the audio and language modalities, we insert a modality adapter composed of Q-Former blocks between the two modules. In our design, we freeze all parameters of the audio model and the LLM, and fine-tune only the modality adapter to learn robust audio–text alignment representations. The fusion model is fine-tuned on triplets in the form $(x^{\text{audio}},\mathbf{p},\mathbf{y})$.
The audio input $x^{\text{audio}}$ is first encoded into a continuous representation by applying Q-Former blocks [24] to multiple intermediate hidden states from the audio encoder. Formally, let $\mathbf{h}^{(\ell)}\in\mathbb{R}^{T\times d}$ denote the output from the $\ell$-th encoder layer, where $T$ represents the temporal dimension and $d$ is the hidden dimension. For each selected layer $\ell$, we define learnable query vectors $\mathbf{Q}^{(\ell)}\in\mathbb{R}^{N\times d}$, where $N$ is the number of queries. The Q-Former computes:
$\mathbf{f}^{(\ell)}=\text{Q-Former}(\mathbf{Q}^{(\ell)},\mathbf{h}^{(\ell)})\in\mathbb{R}^{N\times d}$
The outputs from multiple layers are aggregated using learnable scalar weights $\alpha^{(\ell)}$ with the constraint $\sum_{\ell}\alpha^{(\ell)}=1$, and subsequently projected through a linear layer to match the LLM embedding dimension $d^{\prime}$:
$\mathbf{F}=\text{Linear}\left(\sum_{\ell}\alpha^{(\ell)}\mathbf{f}^{(\ell)}\right)\in\mathbb{R}^{N\times d^{\prime}}$
Optionally, a linguistic representation may be incorporated by transcribing the input audio $x^{\text{audio}}$ using the audio decoder to obtain a text sequence $\mathbf{t}\in\mathbb{R}^{L}$, where $L$ denotes the sequence length. This transcription is then passed through the LLM’s token embedding layer:
$\mathbf{E}=\text{Embed}(\mathbf{t})\in\mathbb{R}^{L\times d^{\prime}}$
When used, the discrete features $\mathbf{E}$ are concatenated with the continuous features $\mathbf{F}$ along the sequence dimension to form the final audio representation:
$\mathbf{A}=[\mathbf{F};\mathbf{E}]\in\mathbb{R}^{(N+L)\times d^{\prime}}$
Otherwise, only the continuous features $\mathbf{F}$ are used.
The resulting audio embeddings $\mathbf{A}$ are passed to the LLM along with the prompt embeddings $\mathbf{P}$ to autoregressively generate the output sequence:
$y_{i}=\text{LLM}(\mathbf{P},\mathbf{A},y_{<i})$
where $y_{i}$ represents the $i$-th token in the generated sequence, and $y_{<i}$ denotes all previously generated tokens. The model undergoes end-to-end optimization using the standard next-token prediction loss computed on the training targets $\mathbf{y}$.
IV Experiment Setup
IV-A Dataset
To construct the training corpus, we collected 50 publicly available datasets spanning a wide range of domains in audio processing. We prioritize datasets containing metadata that cover a comprehensive range of audio information, including paralinguistic features (e.g., pitch, loudness, speaking rate, prosody, timbre, emotional tone, and speaking style), speaker identity attributes (e.g., accent, gender, and age), audio quality indicators (e.g., background noise levels, reverberation, and spoofed or synthetic audio), and environmental or contextual sounds (e.g., animal vocalizations, human actions, ambient sounds, musical instruments, music genres, and natural environments). While the majority of the data are in English, some datasets include Mandarin or code-switching audio samples.
In total, the dataset comprises approximately 7,000 hours of audio: 5,400 hours of speech, 1,000 hours of environmental sounds, and 500 hours of music. For a detailed breakdown of the metadata used in each dataset, we refer readers to our project page for details11footnotemark: 1. The datasets used in this study are categorized as follows:
Speech domain
IEMOCAP [62], DailyTalk [63], GLOBE [64], VCTK-corpus [65], MELD [66], PromptTTS [67], Expresso [68], AccentDB [69], VoxCeleb1 [70], Anispeech [71], MUSAN [72], MSP-IMPROV [73], Fair-speech [74], CREMA-D [75], CAFE [76], EMOVO [77], Speech accent archive [78], EMNS [79], KeSpeech [80], ESD [81], LibriSpeech-c [82], L2Arctic [83], CommonVoice (English and Chinese)[84], EmoV-DB [85], LibriTTS-R [86], Dusha [87], MSP-PODCAST [88], AliMeeting [89], CSZS [90], NTUML2021 [91], Speech Command [92], Libricount [93], Voxlingual [94], ASVspoof [95], BIIC-Podcast [96], and CodecFake  [97, 98].
Environmental sound domain
Audioset [99], AudioCaps [100], Wavcaps [44], Clotho [101], VocalSound [102], UrbanSound8k [103], ESC50 [104], FSD50K [105] , and THMINT-QI [106].
Music domain
Nsynth [107], OpenSinger [108], FMA [109], GTZAN [110] , and Mridangam [111].
Regarding the instruction pool, we curated 4,000 prompts for the speech category and 3,000 prompts for the environmental sound and music categories. Due to the heterogeneity and varying sizes of the source datasets, we applied an upsampling strategy to balance across domains. Each audio sample is paired with multiple prompts, formatted as:
$\{x^{text}\}\{p\}$
All responses were generated using the vLLM toolkit [112], with decoding parameters set to a temperature of 0.05 and a top-p value of 1.0. This process yielded a large-scale dataset of approximately 5 million audio–prompt–response triplets, which we refer to as DeSTA-AQA5M. This dataset serves as the training corpus for DeSTA2.5-Audio. In this work, while we primarily study the performance of Llama3.1-8B-Instruct  [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct [4], the same data construction process can be easily generated by other LLMs within one day on a single NVIDIA A100-80GB GPU.
IV-B Model Specification and Training Setup
We adopt Llama3.1-8B-Instruct and Whisper-large-v3444openai/whisper-large-v3 as the foundational components. We employ a six-layer Q-former architecture [24] with 64 queries as the modality adapter. The query vectors attend to intermediate hidden states from Whisper encoder layers 8, 16, 24, and 32, allowing the model to capture multi-scale acoustic features [113, 114]. For the optional linguistic representation, we provide offline transcriptions for datasets originating from the speech domain. For datasets in the audio and music domains, no transcription is used, and only continuous embeddings derived from the Q-Former are utilized during training. During inference, we employ a lightweight pre-trained voice activity detection (VAD) model [115] to identify the presence of human speech in audio inputs and conditionally activate the Whisper decoder when necessary.
The model implementation is based on the Transformers library [116]. It consists of 8.8 billion total parameters, with 131 million trainable parameters. Training is performed for five epochs using the Adam optimizer [117], a cosine annealing learning rate schedule, and 2000 warm-up steps. The training is conducted on a cluster of 8 NVIDIA A100-80GB GPUs, with a global batch size of 96 and an initial learning rate of 1e-4. The total number of training steps is approximately 250,000.
IV-C Comparison Study
To assess the robustness and effectiveness of our proposed framework, we construct a controlled subset of 500,000 samples from DeSTA-AQA5M. This subset enables systematic analysis of how various training target generation alternatives and model configurations affect model performance. Unless otherwise specified, all comparison experiments are trained for ten epochs using identical model architectures and hyperparameter settings.
In the dataset-level comparison, we fix both the audio descriptions and prompts, and vary only the LLM responsible for generating the training targets. We consider Llama3.1-8B-Instruct [4], Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct [6], Gemma3-12B-it [^6] https://huggingface.co/google/gemma-3-12b-it [7], and Llama3.1-70B-Instruct [^7] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct [4]. This setup facilitates an analysis of how variations in the output distribution across different LLMs influence downstream training outcomes under otherwise controlled conditions. Additionally, to isolate the impact of prompt diversity, we construct a caption-style dataset using a single fixed prompt (i.e., “What can you hear from the audio?”) as proposed in our prior work [9]. For the model-level comparison, we augment the LLM backbone with LoRA adapters [37] of rank 8, applied to the query, key, and value projections of each transformer layer. This increases the number of trainable parameters to 140 million. Additionally, we include a five-epoch variant in which only the number of updates and the learning rate scheduler differ from the ten-epoch setting.
V Evaluation Setup
V-A Benchmarks
To evaluate the capabilities of LALM across instruction-following, perceptual understanding, and reasoning, we adopt a diverse suite of benchmarks. These benchmarks are designed to target distinct challenges in audio-language learning and, when combined, offer a comprehensive evaluation of model behavior across task types, domains, and reasoning levels.
Dynamic-SUPERB Phase-1 [32] evaluates instruction-following and speech understanding across 48 classification tasks. These tasks are categorized into five groups: content (CON), semantic (SEM), paralinguistic (PAR), degradation (DEG), and speaker (SPK). Performance is measured by classification accuracy against ground truth labels.
Dynamic-SUPERB Phase-2 [33] extends the benchmark to 180 tasks by incorporating new contributions from the research community, including regression and open-ended generation tasks across the speech, environmental sound, and music domains. Performance is assessed using task-specific metrics, following the official evaluation guidelines.
MMAU [28] is a benchmark for evaluating advanced audio-language understanding and reasoning across speech, environmental sounds, and music, using multiple-choice question formats. Some questions require expert-level domain knowledge for correct interpretation. We use the test-mini subset, and the performance is evaluated by accuracy against ground truth answers.
SAKURA [29] is a benchmark developed to evaluate single-hop and multi-hop reasoning in LALMs, referred to as SAKURA-Single and SAKURA-Multi, respectively. The benchmark consists of multiple-choice questions. Specifically, single-hop questions assess basic auditory perception, such as identifying sounds like “What animal is making this sound?” In contrast, multi-hop questions require combining auditory cues with external world knowledge and reasoning beyond the immediate input. For example, a multi-hop question might ask, “Is the sound source a mammal or not?”, which requires both recognizing the sound and knowing which animals are classified as mammals. These benchmarks are particularly useful for assessing the internal knowledge of LALM.
Speech-IFEval [36] is a diagnostic benchmark designed to assess whether LALMs retain their instruction-following after cross-modal alignment. Unlike benchmarks that emphasize task-level accuracy, Speech-IFEval introduces the instruction-following rate (IFrate). This metric evaluates whether the model follows output constraints (e.g., limiting response length or generating valid JSON), independent of correct audio interpretation. To quantify degradation in instruction-following, Speech-IFEval further defines the forgetting rate ($\Delta$) as the relative drop in IFrate between an LALM and its backbone LLM. A negative score indicates that the LALM has lost instruction-following ability compared to its text-only backbone.
VoiceBench [30] comprises a collection of tasks designed to evaluate spoken interaction performance. It utilizes text-to-speech (TTS) systems to convert textual instructions into audio inputs, simulating realistic voice-based scenarios. Most tasks in VoiceBench are adapted from existing text-based benchmarks and are primarily content-oriented. We evaluate the performance in accordance with the original guidelines.
V-B Cascade Baselines
We construct two cascade baselines that operate on text-only inputs. The first baseline, ASR+LLM, uses Whisper-large-v3 to transcribe the audio signal into text, which is then directly fed into Llama3.1-8B-Instruct. While this approach omits non-verbal acoustic information, ASR+LLM serves as a strong baseline across multiple benchmarks due to its complete preservation of the LLM’s pre-trained textual knowledge and reasoning ability. To incorporate audio-based information beyond linguistic content, we introduce an extended variant that includes automatic audio captioning (AAC), denoted as ASR+AAC+LLM. Specifically, we use DeSTA2.5-Audio to generate a detailed caption of the audio, which is then concatenated with the ASR transcription to form a richer textual input. This allows the model to reason more effectively about context, emotion, and acoustic scenes, potentially enhancing its performance on tasks that rely on paralinguistic and contextual cues.
V-C Model Inference Setup
While recent studies have demonstrated that careful prompt design, such as chain-of-thought prompting [118], can significantly improve multi-modal performance [119, 120], our goal is to avoid confounding effects caused by prompt engineering. Therefore, we intentionally keep the setting simple throughout this study, and we leave the exploration of more advanced prompting techniques as a direction for future work.
We utilize a set of standardized system prompts to guide the model’s behavior across different evaluation scenarios. The default prompt is “Focus on the audio clips and instruction,” which encourages the model to attend to both modalities and generate informative responses. For benchmarks involving multiple-choice formats, such as Dynamic-SUPERB Phase-1, MMAU, and SAKURA, we append the instruction “Choose one of the options without any explanation” to ensure unambiguous predictions. In VoiceBench, where the instruction is spoken, we adopt the prompt “You are a voice assistant. Act as if you are a natural language partner” to simulate more interactive behavior. For tasks evaluated with word-level metrics, such as ASR tasks using word error rate, we specify “Respond directly without any other words” to suppress auxiliary phrases and ensure fair assessment. We use greedy decoding across all evaluations.
VI Main Results
TABLE I: Evaluation results across Dynamic-SUPERB Phase-1, MMAU, SAKURA and Speech-IFEval. $\dagger$ denotes the cascade reference system result from [36].
Model Dynamic-SUPERB Phase-1 ($\uparrow$) MMAU ($\uparrow$) SAKURA ($\uparrow$) Speech-IFEval ($\uparrow$)
CON SEM PAR DEG SPK Avg Speech Sound Music Avg Single Multi IFrate $\Delta$
ASR+LLM 73.05 59.67 40.50 43.13 43.50 51.71 57.96 38.44 46.41 47.60 40.20 37.40 $\text{93.52}^{\dagger}$
ASR+AAC+LLM 82.77 68.58 45.93 54.87 48.10 60.97 57.96 48.35 47.01 51.10 59.45 51.30
LTU-AS [10] 43.95 36.00 17.14 37.53 40.20 36.11 23.35 9.10 20.60 17.68 40.90 18.10 29.19 -54.90
SALMONN [12] 52.00 50.75 24.50 28.16 33.20 36.44 34.80 41.00 25.50 33.70 42.50 35.30 36.89 -50.20
Qwen-Audio [16] 61.77 47.17 28.64 30.95 41.40 40.79 30.03 44.00 55.25 43.10 73.20 46.90 32.98
Qwen2-Audio-Instruct [17] 77.64 59.17 29.21 43.58 47.90 51.69 42.04 54.95 50.98 49.20 81.20 49.10 47.11
DeSTA2 [9] 79.41 59.42 43.14 51.63 42.50 56.78 55.86 47.75 47.60 50.40 63.00 57.10 89.23 -3.57
DeSTA2.5-Audio 88.95 66.00 63.71 65.08 56.10 69.53 59.16 60.66 52.69 57.50 76.65 69.85 93.89 +0.40
VI-A Results on Dynamic-SUPERB Phase-1, MMAU, SAKURA and Speech-IFEval
Table I presents results across benchmarks, compared to several representative models. To contextualize the benchmark performance, we first consider tasks related to auditory perception, including Dynamic-SUPERB Phase-1 and MMAU, which evaluate models on audio understanding. The SAKURA benchmark, which spans both single-hop and multi-hop reasoning, serves as a bridge between basic perception and higher-level reasoning. Finally, we assess general knowledge and instruction-following through benchmarks like SAKURA-Multi and Speech-IFEval.
VI-A1 Overall Trend
The results reveal a consistent trend in model rankings across multiple benchmarks. Notably, DeSTA2.5-Audio demonstrates consistently superior performance, emerging as the best-performing model with strong generalization across diverse audio-language tasks. It achieves the highest scores on Dynamic-SUPERB Phase-1 (69.53), MMAU (57.50), SAKURA-Multi (69.85), and Speech-IFEval (93.89), indicating its robustness and generalization across multiple domains and conditions.
Compared to DeSTA2, which focused primarily on speech data, DeSTA2.5-Audio shows improved performance attributed to the broader coverage of environmental and musical audios in addition to speech. This demonstrates that our proposed data construction strategy is robust and can be effectively extended to other domains. Interestingly, while Qwen2-Audio-Instruct leveraged a massive 510,000 hours of training data, DeSTA2.5-Audio achieves competitive or superior performance using only 7,000 hours. This highlights the efficiency and effectiveness of our training methodology, which enables strong results with significantly less data.
The cascade baselines (ASR+LLM and ASR+AAC+LLM), while limited to text-based inputs such as transcriptions or audio captions, still perform surprisingly well. In some benchmarks, they even outperform earlier LALMs, such as LTU-AS, SALMONN, and Qwen-Audio. This suggests that an LLM can effectively interpret audio concepts when represented in textual form. Although some fine-grained acoustic information is inevitably lost, these baselines show strong instruction-following ability. In contrast, DeSTA2.5-Audio adopts continuous representations of audio, enabling the model to capture richer and more nuanced information beyond what text alone could convey. During the alignment process, the fusion model learns to interact directly with these audio representations, resulting in more grounded and fine-grained perception. This approach allows DeSTA2.5-Audio to outperform cascade baselines, especially on tasks that require deeper audio understanding.
VI-A2 Task Performance
DeSTA2.5-Audio consistently achieves best performance across a variety subdomains. Within speech-related domains, the model demonstrates strong capabilities not only in content and semantic comprehension but also in more advanced perception, such as paralinguistics, degraded speech, and speaker identity in the Dynamic-SUPERB phase-1. Beyond speech, DeSTA2.5-Audio also performs competitively in the audio and music tracks of the MMAU benchmark, underscoring its versatility in processing diverse auditory inputs.
In the single-hop setting (SAKURA-Single), which requires grounding questions to specific audio segments, Qwen2-Audio-Instruct outperforms DeSTA2.5-Audio (81.20 vs. 76.65). This is likely attributable to Qwen2-Audio-Instruct’s extensive instruction tuning, which enhances its ability to respond to direct queries. However, the performance dynamics shift dramatically in the multi-hop setting (SAKURA-Multi). Here, DeSTA2.5-Audio emerges as the top performer with a score of 69.85, while other models suffer significant drops. For example, Qwen2-Audio-Instruct declines by over 30 points (from 81.20 to 49.10), and LTU-AS drops from 40.90 to 18.10, underscoring their struggles with compositional reasoning. A similar trend can be observed in the Speech-IFEval benchmark. For instance, models such as LTU-AS and SALMONN exhibit low IFrate and high forgetting rates, indicating a significant degradation in both text-based knowledge and instruction-following capabilities. In contrast, DeSTA2.5-Audio and DeSTA2 demonstrate remarkable robustness on the Speech-IFEval benchmark. Their performance remains comparable to or even surpasses the text-based reference systems [36]. The results show that our proposed approach yields a significant advantage in multi-hop reasoning and instruction-following performance, while also delivering strong results on core auditory perception tasks.
Figure 3: LALM performance on the Dynamic-SUPERB Phase-2 benchmark [33], relative to the ASR+LLM baseline. Each cell shows a model’s relative score in a specific domain, with darker blue indicating a higher performance rank. The bottom rows summarize each model’s domain win count (positive scores) and average performance across all domains.
TABLE II: Evaluation results on VoiceBench. Performance is measured using task-specific metrics, where higher scores indicate better performance.
Model AlpacaEval CommonEval WildVoice SD-QA MMSU OpenbookQA BBH IFEval AdvBench Overall
ASR+LLM 4.53 4.04 4.16 70.43 62.43 72.53 69.70 69.53 98.08 77.48
GPT-4o-Audio [1] 4.78 4.49 4.58 75.50 80.25 89.23 84.10 76.02 98.65 86.75
Moshi [47] 2.01 1.60 1.30 15.64 24.04 25.93 47.40 10.12 44.23 29.51
GLM-4-Voice [48] 3.97 3.42 3.18 36.98 39.75 53.41 52.80 25.92 88.08 56.48
DiVA [21] 3.67 3.54 3.74 57.05 25.76 25.49 51.80 39.15 98.27 57.39
Qwen2-Audio-Instruct [17] 3.74 3.43 3.01 35.71 35.72 49.45 54.7 26.33 96.73 55.80
Phi-4 [18] 3.81 3.82 3.56 39.78 42.19 65.93 61.8 45.35 100.00 64.32
VITA-1.5 [46] 4.21 3.66 3.48 38.88 52.15 71.65 55.30 38.14 97.69 64.53
DeSTA2 [9] 4.36 3.33 3.42 56.06 66.43 61.98 66.2 52.63 98.27 69.31
DeSTA2.5-Audio 4.41 3.99 3.80 60.04 60.87 74.06 67.00 66.42 98.27 74.52
VI-B Results on Dynamic-SUPERB Phase-2
Figure 3 presents results on the Dynamic-SUPERB Phase-2 benchmark, which comprises 180 diverse, crowd-sourced tasks, each evaluated using task-specific metrics. To support intuitive visualization and comparison, the benchmark organizes tasks into domains and reports relative scores with respect to an ASR+LLM cascade baseline, which serves as the 0% reference point. This framing emphasizes model capabilities beyond those of a basic cascade pipeline.
Compared to other LALMs, DeSTA2.5-Audio achieves the highest overall performance, ranking first in both win count (14 domains) and average relative score (3.42). These results highlight the model’s robustness and versatility across a wide range of general audio tasks. Notably, DeSTA2.5-Audio does not rely on task-specific instruction tuning. Instead, we leverage diverse and descriptive audio–text pairs, allowing the model to develop a multifaceted understanding of audio while keeping its language ability.
A closer analysis reveals that in certain domains, such as “Quality Assessment, Speech, Voice, Hearing Disorder”, and “Phonetics, Phonology, and Prosody”, our model underperforms the cascade baseline, as indicated by negative relative scores. This suggests that the audio attributes related to addressing these types of tasks are currently underrepresented in our training corpus, thereby limiting generalization performance in these areas. However, the relative scores remain relative to the ASR+LLM baseline, implying marginal deviation despite underperformance. In contrast, some systems exhibit severe performance degradation, for example, LTU-AS in Phonetics, Phonology, and Prosody, and SALMONN in Signal Characteristics Analysis.
Interestingly, rather than producing hallucinated or misleading responses in unfamiliar tasks, DeSTA2.5-Audio often outputs responses like “I don’t have enough information,” demonstrating a form of uncertainty awareness. We consider this behavior particularly valuable for real-world deployment scenarios, where the ability to recognize and communicate uncertainty is critical for trustworthiness and reliability. This trait can be attributed to both the inherent behavior of the backbone LLM and our training strategy, which implicitly encourages such behavior.
TABLE III: Evaluation results across various configurations on representative benchmarks. “Backbone” and “Dataset” refer to the LLM used for LALM training and for dataset generation, respectively. The label “(1-P)” indicates that the dataset was generated using a single descriptive prompt[9]. “PPL” denotes the perplexity of the training targets as computed by the corresponding backbone LLM. “N/A” indicates that the result is unavailable due to degeneration.
ID Backbone Dataset PPL Dynamic-SUPERB Phase-1 MMAU-test-mini SAKURA Speech-IFEval
CON SEM PAR DEG SPK Avg Speech Sound Music Avg Single Multi IFrate $\Delta$
Dataset configuration
Self-generation
A1 Llama3.1 Llama3.1 1.61 89.91 70.50 59.50 56.89 52.70 66.10 57.36 57.66 48.20 54.40 72.45 61.25 91.51 -2.15
A2 Qwen2.5 Qwen2.5 1.27 87.45 65.58 46.86 72.42 59.50 69.94 54.05 65.77 52.10 57.30 78.30 68.65 88.49 -0.24
A3 Llama3.1 Llama3.1 (1-P) 1.58 79.73 60.33 54.57 55.95 55.60 61.71 51.95 56.76 41.62 50.10 55.95 49.85 90.09 -3.67
Cross-model
B1 Llama3.1 Qwen2.5 3.54 N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
B2 Llama3.1 Gemma3-12B 4.85 72.55 70.33 39.86 54.26 52.50 58.18 55.86 48.65 38.02 47.50 52.70 32.25 90.92 -2.79
B3 Llama3.1 Llama3.1-70B 2.20 93.77 73.25 54.64 62.74 57.50 69.44 54.95 56.46 44.61 52.00 75.70 67.35 89.73 -4.06
Model configuration
With LoRA Adapter
C1 Llama3.1-LoRA Llama3.1 1.61 83.00 66.00 57.50 61.82 61.40 66.52 60.06 53.45 52.99 55.50 71.50 64.85 91.17 -2.52
C2 Llama3.1-LoRA Qwen2.5 3.54 87.36 72.92 45.43 70.82 54.10 69.43 59.16 61.26 46.41 55.60 58.75 52.85 84.07 -10.11
With 5 epochs
D1 Llama3.1 Llama3.1 1.61 90.77 70.42 57.50 58.08 53.60 66.56 55.26 57.66 48.20 53.70 70.25 58.80 92.94 -0.62
D2 Llama3.1 Llama3.1 (1-P) 1.58 77.55 60.58 50.57 42.00 40.00 53.51 54.95 54.05 40.12 49.70 58.15 51.70 90.21 -3.54
D3 Llama3.1 Qwen2.5 3.54 74.82 62.67 53.21 53.45 57.70 59.91 57.06 55.56 43.41 52.00 44.30 53.40 91.80 -1.85
VI-C Results on VoiceBench
Table II presents the results on VoiceBench, where the audio input consists of spoken instructions, such as questions or commands, and the model is expected to generate appropriate responses. Unlike previous benchmarks that focus mainly on auditory understanding or instruction-following, VoiceBench emphasizes general knowledge and interactive behavior. To align with these goals, we guide DeSTA2.5-Audio using a system prompt that encourages more natural and conversational responses.
In terms of overall performance, the ASR+LLM pipeline proves to be a very strong baseline, achieving an overall score of 77.48 and outperforming most fusion-based approaches. This indicates that the tasks in VoiceBench are primarily content-driven and can be effectively addressed by accurate transcription paired with a powerful LLM. For comparison, OpenAI’s proprietary GPT-4o-Audio achieves the highest score of 86.75 on VoiceBench, establishing the current state-of-the-art in the field. Despite the simple of cascade systems, most open-source fusion models are still unable to exceed this baseline. This highlights a notable research gap in the field of recent LALM development. These models differ considerably in how they represent audio input. For example, Moshi [47] and GLM-4-Voice [48] adopt unit-based representations, while DiVA [21], Qwen2-Audio-Instruct [17], Phi-4 [18], VITA-1.5 [46] use continuous representations. Furthermore, several models such as Moshi, GLM-4-Voice, and VITA-1.5 are specifically designed for speech-to-speech interaction, making their training procedure more difficult than other speech-to-text model. Among all fusion-based models evaluated, DeSTA2.5-Audio achieves an overall score of 74.52, outperforming models such as VITA-1.5 (64.53) and Qwen2-Audio-Instruct (55.80). It demonstrates balanced and consistent results across a broad range of subtasks, including AlpacaEval, CommonEval, and OpenbookQA. This performance is largely attributed to the strong language ability of DeSTA2.5-Audio, where it not only follows the system prompt to adapt its behavior into an interactive style, but also responds appropriately based on its inherent knowledge.
In summary, the proposed DeSTA2.5-Audio model demonstrates state-of-the-art performance or comparable across a wide range of audio-language benchmarks. It outperforms both previous versions (e.g., DeSTA2) and competing models on key tasks such as Dynamic-SUPERB Phase-1, MMAU, SAKURA-Multi, and Speech-IFEval (Section VI-A). Notably, DeSTA2.5-Audio achieves competitive results with significantly less training data (7,000 hours) compared to models like Qwen2-Audio-Instruct (510,000 hours). It also exhibits robustness in knowledge-intensive tasks and shows strong instruction-following ability. Evaluations on Dynamic-SUPERB Phase-2 (Section VI-B) and VoiceBench (Section VI-C) further validate its generalization and robustness, highlighting its potential as a general-purpose LALM.
VII Comparison Studies
As shown in Table III, we evaluate the proposed self-generation approach under controlled experimental setups using identical audio-description pairs from a 500K subset of DeSTA-AQA5M. The results are analyzed across four settings to provide a comprehensive understanding. The self-generation setting represents the proposed configuration, where the same backbone LLM is used both to construct the training target and to serve as the LALM backbone. In this case, there is no distribution mismatch between the model and the training data. In the cross-model setting, different LLMs are used to generate the training targets. This setup simulates a distribution mismatch, such as when human annotators or alternative LLMs are employed to create the training corpus. This setup is widely adopted in previous LALM studies [8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Building upon these two setups, we explore configurations with LoRA adapters integrated into the backbone LLM. This allows us to examine whether introducing additional trainable parameters helps mitigate performance degradation due to data mismatch or enhances generalization. Lastly, to examine the effect of training duration on model performance, we adopt a 5-epoch setting in contrast to our default 10-epoch setup. This allows us to investigate how increased exposure to the training data influences model performance. To quantify the degree of distribution mismatch, we report the perplexity (PPL) of the training targets as computed by the corresponding backbone LLM. A lower perplexity indicates the model is more familiar with the training target, suggesting less distributional discrepancy.
VII-A Comparison between Self-generated and Cross-model Settings
As indicated in Table III, the self-generated training data exhibit consistently lower perplexity, suggesting that the generated responses align well with the distribution of the backbone LLM. When comparing Llama3.1 (A1) and Qwen2.5 (A2), Qwen2.5 consistently outperforms Llama3.1 across all benchmarks. This performance gap may be attributed to Qwen2.5’s stronger text generation capabilities. While Qwen2.5’s performance in basic content comprehension tasks is relatively on par with Llama3.1, it demonstrates better performance in other areas, such as speaker category in Dynamic-SUPERB Phase-1, and environmental sound and music understanding in MMAU. Previous evaluations on text-based benchmarks have also shown that Qwen2.5 exhibits superior reasoning and mathematical abilities compared to Llama3.1 [121]. However, there is currently no conclusive evidence indicating a corresponding advantage in auditory perception, which warrants further investigation. Nevertheless, under identical training conditions, our experimental results suggest that Qwen2.5 serves as a more effective backbone LLM than Llama3.1. These findings also indicate that our training framework generalizes well across different LLMs.
Prompt diversity also plays a significant role in model performance, particularly in the comparison between A1 and A3. In A3, we adopt the self-generation setup using a single descriptive prompt (1-p), as described in our previous work [9], which already demonstrates strong zero-shot generalization. By simply increasing prompt diversity, as done in A1, we further enrich the training target and enhance the overall effectiveness of the training method. Notably, these results are achieved without the need for any task-specific instruction pairs. This highlights the strength of the self-generation design. Even when data construction relies solely on randomly sampled prompts, the model can still achieve zero-shot generation by leveraging the inherent capabilities of the LLM.
When comparing self-generation and cross-model settings, training targets in the cross-model setting result in higher perplexity, suggesting that the backbone LLM is less familiar with the data distribution. For example, while training Qwen2.5 on Qwen2.5-generated data (A2) produces strong results, training Llama3.1 on Qwen2.5-generated data (B1) leads to model degeneration [122], with outputs containing repetitive or nonsensical tokens. Similarly, training Llama3.1 on data generated by Gemma3-12B (B2) fails to match the performance observed in the self-generation setting (A1). These results support our distribution mismatch hypothesis and emphasize the importance of using a self-generated configuration, even when the annotator LLM is more capable. We also explore using Llama3.1-70B to generate training data (B3), representing a more powerful model from the same family. In this case, the lower perplexity (2.20) suggests a closer alignment between the training data and Llama3.1’s distribution. However, compared to A1, B3 achieves better performance on Dynamic-SUPERB and SAKURA, but underperforms on MMAU and Speech-IFEval. This indicates that using a stronger model does not necessarily lead to consistent improvements across all tasks.
VII-B With LoRA Adapter
In the LoRA adapter setting, we introduce trainable parameters to the backbone LLM, which is expected to increase model capacity and help mitigate the distribution mismatch problem. In the self-generation setup (C1), where the dataset is well-aligned with the backbone LLM, we find that adding LoRA layers yields similar or only slightly better performance. This indicates that, under self-generation settings, incorporating LoRA adapters does not provide significant advantages. In other words, fine-tuning a lightweight modality adapter is sufficient for cross-modal alignment when using our proposed training framework, where the model focuses on learning auditory concepts without being hindered by stylistic or distributional mismatches. Interestingly, when training with Qwen2.5-generated data (C2), the performances in audio processing benchmarks, such as Dynamic-SUPERB, MMAU, and SAKURA-Single, are comparable to the self-generation setup (A2). However, they experience a significant degradation in SAKURA-Multi and Speech-IFEval, which requires additional text knowledge and instruction-following ability. This difference indicates that while adding a LoRA adapter can help mitigate the distribution mismatch and perform well on in-domain tasks, it may still degrade the model’s general capabilities when evaluated on benchmarks that require knowledge from LLM pretraining. This reveals a critical design limitation in current LALM training strategies. Models such as LTU-AS and SALMONN attempt to address the catastrophic forgetting problem by introducing LoRA adapter layers to the LLM. However, our experimental results suggest that reducing the discrepancy between training data and model distribution is a more critical factor for preserving generalization ability than architectural modifications alone.
VII-C With 5 epochs
With the 5-epoch setting, we investigate the impact of training duration on model performance. Consistent with discussion in Section VII-A, the 5-epoch results (D1 and D2) show that prompt diversity not only improves effectiveness but also enhances training efficiency. Despite being trained for only half the number of epochs, the models achieve performance comparable to their 10-epoch counterparts (A1). Notably, while D2 continues to improve with extended training (as in A3), convergence is slower and the final performance remains worse, indicating that diverse training target is also essential for achieving better alignment. In contrast, although D3 achieves non-trivial performance with only 5 epochs of training, B1 demonstrates that extended training under distribution mismatch leads to model degeneration. These findings underscore our key motivation: effective cross-modal alignment requires repeated exposure to align audio representations across epochs. When the training data is matched with the backbone model, performance improves steadily without degrading the model’s inherent language capabilities. In contrast, learning from mismatched data places a heavier burden on the model, ultimately resulting in suboptimal performance and forgetting of its pretrained linguistic abilities.
VIII Conclusion
In this work, we introduced DeSTA2.5-Audio, a general-purpose LALM built on the self-generated cross-modal alignment framework, DeSTA. By allowing the backbone LLM to generate its own training targets, DeSTA maintains distributional consistency and mitigates the catastrophic forgetting problem commonly observed in prior works. Under this framework, we successfully develop a robust and generalizable LALM without relying on task-specific instruction data. We constructed DeSTA-AQA5M, a large-scale, task-agnostic dataset comprising 5 million audio-text pairs from 50 diverse audio datasets. Trained solely on this corpus, DeSTA2.5-Audio achieves state-of-the-art or competitive results across benchmarks such as Dynamic-SUPERB (Phase-1 and Phase-2), MMAU, SAKURA, Speech-IFEval, and VoiceBench. Our findings highlight the importance of training data quality over quantity in LALM development. Compared to previous studies, our self-generation strategy offers a scalable and robust solution with better generalization.
While our current design leverages text description as an intermediate bridge, we acknowledge that not all acoustic nuances can be effectively captured through textual representations. In future work, we will continue to explore methods that enable LALMs to better capture subtle, non-text-expressible audio features.
IX Acknowledgment
The authors thank NVIDIA Taiwan AI R&D Center for the TRDC budget support and contributions to this research. This work has also been supported by NVIDIA Academic Grant from the NVIDIA Developer Program 2025.